OpenStudy (anonymous):

Simplify This Expression (with steps): (x^2-4)/(2x-4) thanks

OpenStudy (anonymous):

I guess you have to do long division of 2x-4 into x^2 - 4

OpenStudy (anonymous):

ugh, why do they make students do this?

OpenStudy (anonymous):

(x^2-4)/2x-4) = [(x+2)(x-2)]/2(x-2) = (x+2)/2

OpenStudy (anonymous):

here is how to do the long division: http://i.imgur.com/GGPzH.gif

OpenStudy (anonymous):

thank you both but the question was asking for simplifying not actually solving

OpenStudy (anonymous):

that's what we did

OpenStudy (anonymous):

the expression is equal to x/2 + 1 (my way ) or equivalently, (x+2) / 2 (his way)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

so did you understand the answers then?

OpenStudy (anonymous):

yeah i got it, many thanks

OpenStudy (anonymous):